if arg(z1/z2)=p/2 then find the value of |(z1+z2)/z1-z2)|

Grade:

2 Answers

jitender lakhanpal
10 years ago

|(z1+z2)/z1-z2)| = 1


|(z1+z2)|= |z1|2   + |z1|+  2|z1||z2|cos (θ1 -θ2)

and
|(z1-z2)|= |z1|2   + |z1|2    - 2|z1||z2|cos (θ1 -θ2)

where  θ1 and θ2   are arguments of z1 and z2


θ1 -θ2 = pie/2       so  cos (θ1 -θ2) = 0

so both of them are equal

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

|(z1+z2)/z1-z2)| = 1
|(z1+z2)|2 = |z1|2 + |z1|2 + 2|z1||z2|cos (θ1 -θ2)
and
|(z1-z2)|2 = |z1|2 + |z1|2 - 2|z1||z2|cos (θ1 -θ2)
where θ1 and θ2 are arguments of z1 and z2
θ1 -θ2 = pie/2 so cos (θ1 -θ2) = 0
so both of them are equal

Thanks and Regards